OpenStudy (anonymous):

Session 9: Matrix Multiplication

OpenStudy (anonymous):

Seems like the answer to problem (i) in the solutions is incorrect. \[\left[\begin{matrix}1 & 3 \\ 4 & 5\end{matrix}\right] *\left[\begin{matrix}1 & 3 \\ 4 & 5\end{matrix}\right]\] = \[\left[\begin{matrix}13 & 8 \\ 24 & 37\end{matrix}\right]\] ...should read... \[\left[\begin{matrix}13 & 18 \\ 24 & 37\end{matrix}\right]\] Is the correct?

OpenStudy (phi):

yes, they probably have a typo your answer is correct.